Given fixed cost:$40,500
Total variable cost: $192,000
Unit selling price:$100
Units sold:15,000
Find: CM
Find break-even units.
Find units required to sell if need $100,000 profit
Answer #1

Total contribution margin = Volume * (Selling price - Unit variable cost)

= Volume * Selling price - Volume * Unit variable cost

= Volume * Selling price - Total variable cost

= 15,000 x 100 - 192,000

= $1,308,000

Unit contribution margin = Total contribtion margin / volume = 1308000 / 15000 = $87.2

Break-even units = Fixed cost / unit contribution margin = 40500 / 87.2 = 465 units

Profit = Volume * Unit contribution margin - Fixed cost

or, 100,000 = Volume * 87.2 - 40500

or, Volume to attain a profit of $100,000 = 140500 / 87.2 = 1612 units